# Linearity of the co-moving velocity

### 21 Feb 2024 v1

@@ 3 Derivation
Hence, $b$ is a constant. Combining this result with equation (16) gives the final result,
$v_{m}(a,v_{p}^{\prime})=a+b\ v_{p}^{\prime}\;.$ (19)
@@
The co-moving velocity $v_{m}$ is linear in $v_{p}^{\prime}$.
@@ 4 Consequences for relative permeability theory
where $\mu_{w}$ and $\mu_{n}$ are viscosities, $K$ the permeability and $k_{rw}(S_{w})$ and $k_{rn}(S_{w})$ are the relative permeabilities. The saturations appear in the expressions since $v_{w}$ and $v_{n}$ are the seepage velocities.
@@
where I have defined the viscosity ratio $M=\mu_{w}/\mu_{n}$. I now combine these two expressions with the co-moving velocity, equation (19) using equation (6). I also make a change in notation by setting $a\to av_{0}$, as in roy2022co . The resulting equation is
@@
A simplified version of this equation was used in Alzubaidi et al. alzubaidi2023impact to determine the non-wetting relative permeability when the wetting permeability, $a=0$ and $b$ is known. One way to use this equation would be to plug the known $k_{rw}$ into it and then find $k_{rn}$ through numerical integration.
It is also possible to find analytic solutions to equation (27). I give an example in the following. Picci and Battiato picchi2019relative suggested the semi-empirical form (note that their viscosity ratio is the inverse of the definition I use)
@@
By solving for the saturation for which $v_{p}^{\prime}=0$, using equations (28) and (29), we find that
$[v_{n}-v_{w}]_{v_{p}^{\prime}=0}=k_{rn}^{0}\;.$ (32)
@@
The Corey empirical relative permeabilities are, on the other hand, not solutions to equation (27). That is, when the Corey exponents are equal, they are solutions, but with an unphysical $a=0$. When the exponents are different, they are not solutions to the equation for any combination of $b$ and $a$.
@@ 5 Discussion and Conclusion
I have in this paper shown that the co-moving velocity is of the form given by equation (19). The derivation is based on two assumptions, 1. the necessity of an additive velocity scale $a$, and 2. the invariance of $v_{m}$ with respect to a change of scale $v_{0}$. Both assumptions are weak and does not hinge on any restrictions on the constitutive law for $v_{p}$, linear or non-linear. In order to connect relative permeability theory to the formalism for which $v_{m}$ is a central variable, the specific constitutive law
$v_{p}(p,S_{w})=\left[k_{rw}(S_{w})+Mk_{rn}(S_{w})\right]v_{0}(p)\;,$ (33)
is used. Equation (6), which is the combination of the relative permeability constitutive laws with the linear equation for $v_{m}$, (19), leads to a relation between the two relative permeabilities: knowing one, gives the other one.
The parameter $a=[v_{n}-v_{w}]_{v_{p}^{\prime}=0}$, equation (11) is understood. However, the second parameter, $b$ is not. It typically takes on a value around 0.6–0.8 and seems quite insensitive to the contact angle between the fluids and the matrix as shown by Alzubaidi et al. alzubaidi2023impact . It is an open question why. Furthermore, from equation (6) one should think that $b=1$. It is not. Why?

### 25 Mar 2024 v2

@@ 4 Dimensional analysis of the co-moving velocity
i.e., a constant. From equation (47) I then have that the function $b(x)$ is also a constant. Combining this result with equation (33) gives the final result,
$v_{m}(a,v_{t}^{\prime})=a+b\ v_{t}^{\prime}\;.$ (49)
@@
Hence, the co-moving velocity $v_{m}$ is linear in $v_{t}^{\prime}$ if and only if $a$ is an additive constant.
The assumption that $a$ is an additive constant makes it possible to identify it. From equation (10) I have that
$a=v_{m}(a,0)=[v_{n}-v_{w}]_{v_{t}^{\prime}=0}\;.$ (50)
This identification is only possible if $v_{t}^{\prime}=0$ falls within the physical range of the parameters; in other words, $v_{t}$ has a minimum in the range $0\leq S_{w}\leq 1$. If there is no such minimum, it is still possible to define $a$, but it will not have the simple physical interpretation given in equation (50).
4.1 The example: co-moving velocity
The co-moving velocity in the example is given in equation (23). In order to calculate $a$ according to equation (46), I set $v_{t}^{\prime}=0$ in the equation (21) giving the corresponding saturation $S_{w}=1/2C$. I use the equations for $v_{w}$ and $v_{n}$, (14) and (15) together with the definition of $C$, equation (19),
$a=[v_{n}-v_{w}]_{v_{t}^{\prime}=0}=v_{n}^{0}-v_{w}^{0}+\Delta v_{n}-\frac{1}{2C}(\Delta v_{w}+\Delta v_{n})=\frac{v_{n}^{0}-v_{w}^{0}}{2}\;.$ (51)
We see that $v_{w}^{0}$ and $v_{n}^{0}$ appear in the expression in spite of equation (38). This is so since switching from $S_{w}$ to $v_{t}^{\prime}$ as control variable reintroduces them.
@@ 5 Consequences for relative permeability theory
where $\mu_{w}$ and $\mu_{n}$ are viscosities, $K$ the permeability and $k_{rw}(S_{w})$ and $k_{rn}(S_{w})$ are the relative permeabilities. The saturations appear in the denominators since $v_{w}$ and $v_{n}$ are seepage velocities.
@@
where I have defined the viscosity ratio $M=\mu_{w}/\mu_{n}$. The average velocity, equation (8), is then given by
$v_{t}(p,S_{w})=\left[k_{rw}(S_{w})+Mk_{rn}(S_{w})\right]v_{0}(p)\;.$ (57)
I now combine equations (55) and (56), with the co-moving velocity, equation (10) using equation (13). I also make a change in notation by setting $a\to av_{0}$, as in Roy et al. roy2022co . The resulting equation is
@@
A simplified version of this equation was used in Alzubaidi et al. alzubaidi2023impact to determine the non-wetting relative permeability when the $k_{rw}$, $a=0$ and $b$ are known.
It is also possible to find analytic solutions to equation (58). I give an example in the following. Picchi and Battiato picchi2019relative suggested the semi-empirical form (note that their viscosity ratio is the inverse of the definition I use)
@@
By solving for the saturation for which $v_{t}^{\prime}=0$, using equations (59) and (60), I find that
$[v_{n}-v_{w}]_{v_{t}^{\prime}=0}=k_{rn}^{0}\;.$ (63)
@@
The Corey empirical relative permeabilities
$k_{rw}=k_{rw}^{0}S_{w}^{\mu_{w}}\;,$ (64)
$k_{rn}=k_{rn}^{0}S_{n}^{\mu_{n}}\;,$ (65)
are, on the other hand, not solutions to equation (58), except when $\mu_{w}=\mu_{n}$, they are solutions, but with $a=0$. In that case I find
$b=1-\frac{1}{\mu_{w}}\;.$ (66)
The example I gave in Section 2.1 was based on the Corey relative permeabilities with exponents $\mu_{w}=\mu_{n}=2$. The non-zero $a$ that was found, equation (25), came from the assumption that $v_{w}^{0}$ and $v_{n}^{0}$ were non-zero. The generalized Darcy equations (52) and (53) assumes $v_{w}^{0}$ and $v_{n}^{0}$ are both zero.
@@ 6 Discussion and Conclusion
I have in this paper shown that the co-moving velocity is of the form given by equation (13) given that two conditions are in place: 1. the average seepage velocity is characterized by two parameters with dimensions of velocity, allowing for additional dimensionless parameters, and 2. the necessity of an additive velocity scale $a$ in the co-moving velocity. Both assumptions are weak and does not hinge on any further restrictions on the constitutive law for $v_{t}$, linear or non-linear. Equation (58), which is the combination of the relative permeability constitutive laws (52) and (53) with the linear equation for $v_{m}$, (10), leads to a relation between the two relative permeabilities: knowing one, gives the other one.
The parameter $a=[v_{n}-v_{w}]_{v_{t}^{\prime}=0}$, equation (28) may be given a straight-forward physical interpretation when $v_{t}=0$ falls within the physical range of the parameters. However, the second parameter, $b$ is not. It typically takes on a value around 0.6–0.8 and seems quite insensitive to the contact angle between the fluids and the matrix as shown by Alzubaidi et al. alzubaidi2023impact . It is an open question why.